OpenStudy (anonymous):

The Mogul Runners ski club planned a trip to Park City. Of the total number of club members, 11 signed up to go. If this is 55% of the club, how many members does the ski club have?

OpenStudy (lucaz):

55% is the same as 55/100 so if 55% = 11 -> 55/100 = 11/x just solve for x to find the n of total members